1/4x is greater than or equal to -5/8. The solution set is??? Type an inequality. use integers or fractions for any numbers in the expresson. Thank you
Answer by checkley79(3341)

1/4x>=-5/8
X/4>=-5/8
X>=-5/8*4
X>=-20/8
X>=-2.5 ANS.
PROOF:
LET X=-2.4
-2.4/4>=-5/8
-.600>=-.625